# On the Embedding Collapse When Scaling up Recommendation Models

### v1

@@ Abstract
Recent advances in deep foundation models have led to a promising trend of developing large recommendation models to leverage vast amounts of available data. However, we experiment to scale up existing recommendation models and observe that the enlarged models do not improve satisfactorily. In this context, we investigate the embedding layers of enlarged models and identify a phenomenon of embedding collapse, which ultimately hinders scalability, wherein the embedding matrix tends to reside in a low-dimensional subspace. Through empirical and theoretical analysis, we demonstrate that the feature interaction module specific to recommendation models has a two-sided effect. On the one hand, the interaction restricts embedding learning when interacting with collapsed embeddings, exacerbating the collapse issue. On the other hand, feature interaction is crucial in mitigating the fitting of spurious features, thereby improving scalability. Based on this analysis, we propose a simple yet effective multi-embedding design incorporating embedding-set-specific interaction modules to capture diverse patterns and reduce collapse. Extensive experiments demonstrate that this proposed design provides consistent scalability for various recommendation models.
@@ 1 Introduction
Recommender systems are significant machine learning scenarios that predict users’ actions on items based on multi-field categorical data (Zhang et al., 2016), which play an indispensable role in our daily lives to help people discover information about their interests and have been adopted in a wide range of online applications, such as E-commerce, social media, news feeds, and music streaming. Recently, researchers have developed deep-learning-based recommendation models to dig feature representations flexibly. These models have been successfully deployed across a multitude of application scenarios, thereby demonstrating their widespread adoption and effectiveness.
In recommender systems, there is a tremendous amount of Internet data, while mainstream models typically tuned with embedding size 10 (Zhu et al., 2022) do not adequately capture the magnitude of the available data. Motivated by the advancement of large foundation models (Kirillov et al., 2023; OpenAI, 2023; Radford et al., 2021; Rombach et al., 2022), which benefit from increasing parameters, it would be a promising trend to scale up the recommendation model size. However, we experiment to increase the embedding size of mainstream recommendation models (Qu et al., 2016; Lian et al., 2018; Wang et al., 2021), as shown in Figure 1a, and find an unsatisfactory improvement or even performance drop. This suggests a deficiency in the scalability of existing architecture designs, constraining the maximum potential for recommender systems.
@@
We take a spectral analysis on the learned embedding matrices based on singular value decomposition and exhibit the normalized singular values in Figure 1b. Surprisingly, most singular values are significantly small, i.e., the learned embedding matrices are nearly low-rank, which we refer to as the embedding collapse phenomenon. With the enlarged model size, the model does not learn to capture a larger dimension of information, implying a learning process with ineffective parameter utilization, which restricts the scalability.
In this work, we study the mechanism behind the embedding collapse phenomenon through empirical and theoretical analysis. We shed light on the two-sided effect of the feature interaction module, the characteristic of recommendation models to model higher-order correlations, on model scalability. On the one hand, interaction with collapsed embeddings will constrain the embedding learning and, thus, in turn, aggravate the collapse issue. On the other hand, the feature interaction also plays a vital role in reducing overfitting when scaling up models.
Based on our analysis, we conclude the principle to mitigate collapse without suppressing feature interaction about how to design scalable models. We propose multi-embedding as a simple yet efficient design for model scaling. Multi-embedding scales the number of independent embedding sets and incorporates embedding-set-specific interaction modules to jointly capture diverse patterns. Our experimental results demonstrate that multi-embedding provides scalability for extensive mainstream models, pointing to a methodology of breaking through the size limit of recommender systems.
@@
  • To the best of our knowledge, we are the first to point out the non-scalability issue for recommendation models and discover the embedding collapse phenomenon, which is an urgent problem to address for model scalability.
  • We shed light on the two-sided effect of the feature interaction process on scalability based on the collapse phenomenon using empirical and theoretical analysis. Specifically, feature interaction leads to collapse while providing essential overfitting reduction.
  • Following our concluded principle to mitigate collapse without suppressing feature interaction, we propose multi-embedding as a simple unified design, which consistently improves scalability for extensive state-of-the-art recommendation models.
@@ 2 Preliminaries
Recommendation models aim to predict an action based on features from various fields. Throughout this paper, we consider the fundamental scenario of recommender systems, in which categorial features and binary outputs are involved. Formally, suppose there are $N$ fields, with the $i$-th field denoted as $\mathcal{X}_{i}=\{1,2,...,D_{i}\}$ where $D_{i}$ denotes the field cardinality. The value of $D_{i}$ may vary in a wide range, adding difficulty to recommender systems. Let
@@ 3 Embedding Collapse
Singular value decompostion has been widely used to measure the collapse phenomenon (Jing et al., 2021). In Figure 1b, we have shown that the learned embedding matrices of recommendation models are approximately low-rank with some extremely small singular values. To determine the degree of collapse for such matrices with low-rank tendencies, we propose information abundance as a generalized quantification.
Definition 1 (Information Abundance)
@@
Figure 2: Visualization of information abundance on the Criteo dataset. The fields are sorted by their cardinalities.
@@ 4 Feature Interaction Revisited
In this section, we delve deeper into the embedding collapse phenomenon for recommendation models. Our investigation revolves around two questions: (1) How is embedding collapse caused? (2) How to properly mitigate embedding collapse for scalability? Through empirical and theoretical studies, we shed light on the two-sided effect of the commonly employed feature interaction module on model scalability.
4.1 Interaction-Collapse Law
To determine how feature interaction leads to embedding collapse, it is inadequate to directly analyze the raw embedding matrices since the learned embedding matrix results from interactions with all other fields, making it difficult to isolate the impact of field-pair-level interaction on embedding learning. To overcome this obstacle, we propose a solution involving experiments on a line of models equipped with sub-embedding modules generated from the raw embeddings and field individually designed for feature interaction. By examining the collapse of all these sub-embeddings, we can effectively discern the interaction effects between the two fields in a disentangled manner and establish a high-level law.
Evidence I: Experiments on FFM.
where sub-embedding ${\bm{E}}_{i}^{\to j}\in\mathbb{R}^{D_{i}\times K/(N-1)}$ is only used when interacting field $i$ with field $j$ for $j\neq i$. To determine the collapse of sub-embedding matrices, we calculate $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ for all $i,j$ and show them in Figure 3a. For convenience, we pre-sort the field indices by the ascending order of information abundance, i.e., $i$ is ordered according to $\mathrm{IA}({\bm{E}}_{i})$, similar to $j$. We can observe that $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ is approximately increasing along $i$, which is trivial since ${\bm{E}}_{i}^{\to j}$ is simply a split of ${\bm{E}}_{i}$. Interestingly, another correlation can be observed that the information abundance of sub-embeddings is co-influenced by the fields it interacts with, reflected by the increasing trend along $j$, especially with larger $i$. This is amazing in the sense that even using independent embeddings to represent the same field features, these embeddings get different information abundance after learning. For instance, we calculate the summation of $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ over $j$ or $i$ to study the effect of the other single variable, shown in Figure 3b and Figure 3c. Both of them show an increasing trend, confirming the co-influence of $i$ and $j$.
Refer to caption (a) $\mathrm{IA}({\bm{E}}_{i}^{\to j})$. Refer to caption (b) $\sum\limits_{j=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$. Refer to caption (c) $\sum\limits_{i=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$.
Refer to caption (d) $\mathrm{IA}({\bm{E}}_{i}^{\to j})$. Refer to caption (e) $\sum\limits_{j=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$. Refer to caption (f) $\sum\limits_{i=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$.
Figure 3: Visualization of information abundance of sub-embedding matrices for FFM (left) and DCNv2 (right), with field indices sorted by information abundance of corresponding raw embedding matrices. Higher or warmer indicates larger. It is observed that $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ are co-influenced by both $\mathrm{IA}({\bm{E}}_{i})$ and $\mathrm{IA}({\bm{E}}_{j})$.
Evidence II: Experiments on DCNv2.
An improved deep & cross network (DCNv2) (Wang et al., 2021) incorporates a crossing network which is parameterized with transformation matrices ${\bm{W}}_{i\to j}$ (Sun et al., 2021) over each field pair to project an embedding vector from field $i$ before interaction with field $j$. By collecting all projected embedding vectors, DCNv2 can be regarded to implicitly generate field-aware sub-embeddings ${\bm{E}}_{i}^{\to 1},{\bm{E}}_{i}^{\to 2},...,{\bm{E}}_{i}^{\to N}$ to interact with all fields from embedding matrix ${\bm{E}}_{i}$, with
@@ 4.1 Interaction-Collapse Law
DCNv2 consists of multiple stacked cross layers, and for simplification, we only discuss the first layer throughout this paper. Similar to Evidence I, we calculate $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ together with the axis-wise summations and show them in the right part of Figure 3. Consistent with previous observation as FFM, the information abundance of sub-embedding matrices approximately increases along $j$ with the same $i$, even though they are projected from the same embedding matrix ${\bm{E}}_{i}$.
Summary: How is collapse caused in real-world scenarios?
In the cases of FFM and DCNv2, the sub-embeddings obtained from the raw embedding, either separated or projected, provide insights into the interactions between different fields. Evidence I&II highlight that the information abundance in a sub-embedding matrix is greatly impacted by the field it interacts with, or specifically, interacting with a field with a low-information-abundance embedding matrix will result in a more collapsed sub-embedding. By further considering the fact that sub-embeddings reflect the effect when fields interact, we conclude an inherent mechanism of feature interaction in recommendation models, leading us to propose the interaction-collapse law from a higher perspective:
{tcolorbox}
[colback=blue!2!white,leftrule=2.5mm,size=title] Finding 1 (Interaction-Collapse Law). In feature interaction of recommendation models, fields with low-information-abundance embeddings constrain the information abundance of other fields, resulting in collapsed embedding matrices.
While the interaction-collapse law is derived from sub-embedding-based models, it is important to note that these sub-embeddings serve as analytical tools that reveal the underlying mechanism of feature interaction. Therefore, this law is applicable to general recommendation models instead of only the models with sub-embeddings. The interaction-collapse law generally suggests that feature interaction is the primary catalyst for collapse, thereby imposing constraints on the ideal scalability.
We now present how collapse is caused by feature interaction in recommendation models from a theoretical view. For simplicity, we consider an FM-style (Rendle, 2010) feature interaction. Formally, the interaction process is defined by
@@
where $h$ is the combined feature as mentioned before. Without loss of generality, we discuss one specific row ${\bm{e}}_{1}$ of ${\bm{E}}_{1}$ and keep other embedding matrices fixed. Consider a minibatch with batch size $B$. Denote $\sigma_{i,k}$ as the $k$-th singular value of ${\bm{E}}_{i}$, similar for ${\bm{u}}_{i,k}$, ${\bm{v}}_{i,k}$. We have
$\frac{\partial\mathcal{L}}{\partial{\bm{e}}_{1}}$ $=\frac{1}{B}\sum_{b=1}^{B}\frac{\partial\ell^{(b)}}{\partial h^{(b)}}\cdot\frac{\partial h^{(b)}}{\partial{\bm{e}}_{1}}=\frac{1}{B}\sum_{b=1}^{B}\frac{\partial\ell^{(b)}}{\partial h^{(b)}}\cdot\sum_{i=2}^{N}{\bm{E}}_{k}^{\top}\bm{1}_{x_{i}^{(b)}}$
@@
$=\sum_{i=2}^{N}\sum_{k=1}^{K}\alpha_{i,k}\sigma_{i,k}{\bm{v}}_{i,k}=\sum_{i=2}^{N}{\bm{\theta}}_{i}$
Figure 4: $\mathrm{IA}({\bm{E}}_{1})$ w.r.t. training iterations for toy experiments. “Small” and “Large” refers to the cardinality of $\mathcal{X}_{3}$.
The equation means that the gradient can be decomposed into field-specific terms. We analyze the component ${\bm{\theta}}_{i}$ for a certain field $i$, which is further decomposed into spectral for the corresponding embedding matrix ${\bm{E}}_{i}$.
From the form of ${\bm{\theta}}_{i}$, it is observed that $\{\alpha_{i,k}\}$ are $\bm{\sigma}_{i}$-agnostic scalars determined by the training data and objective function. Thus, the variety of $\bm{\sigma}_{i}$ significantly influences the composition of ${\bm{\theta}}_{i}$. For those larger $\sigma_{i,k}$, the gradient component ${\bm{\theta}}_{i}$ will be weighted more heavily along the corresponding spectral ${\bm{v}}_{i,k}$. When ${\bm{E}}_{i}$ is low-information-abundance, the components of ${\bm{\theta}}_{i}$ weigh imbalancely, resulting in the degeneration of ${\bm{e}}_{1}$. Since different ${\bm{e}}_{1}$ affects only $\alpha_{i,k}$ instead of $\sigma_{i,k}$ and ${\bm{v}}_{i,k}$, all rows of ${\bm{E}}_{1}$ degenerates in similar manners and finally form a collapsed matrix.
To further illustrate, we conduct a toy experiment over synthetic data. Suppose there are $N=3$ fields, and we set $D_{3}$ to different values with $D_{3}<K$ and $D_{3}\gg K$ to simulate low-information-abundance and high-information-abundance cases, which matches the diverse range of the field cardinality in real-world scenarios. We train ${\bm{E}}_{1}$ while keeping ${\bm{E}}_{2},{\bm{E}}_{3}$ fixed. Details of experiment setups are discussed in Appendix A. We show the information abundance of ${\bm{E}}_{1}$ along the training process for the two cases in Figure 4. It is observed that interacting with a low-information-abundance matrix will result in a collapsed embedding matrix.

### v2

@@ Abstract
Recent advances in foundation models have led to a promising trend of developing large recommendation models to leverage vast amounts of available data. Still, mainstream models remain embarrassingly small in size and naïve enlarging does not lead to sufficient performance gain, suggesting a deficiency in the model scalability. In this paper, we identify the embedding collapse phenomenon as the inhibition of scalability, wherein the embedding matrix tends to occupy a low-dimensional subspace. Through empirical and theoretical analysis, we demonstrate a two-sided effect of feature interaction specific to recommendation models. On the one hand, interacting with collapsed embeddings restricts embedding learning and exacerbates the collapse issue. On the other hand, interaction is crucial in mitigating the fitting of spurious features as a scalability guarantee. Based on our analysis, we propose a simple yet effective multi-embedding design incorporating embedding-set-specific interaction modules to learn embedding sets with large diversity and thus reduce collapse. Extensive experiments demonstrate that this proposed design provides consistent scalability and effective collapse mitigation for various recommendation models. Code is available at this repository: https://github.com/thuml/Multi-Embedding.
Machine Learning, ICML
@@ 1 Introduction
Recommender systems are important machine learning scenarios that predict users’ actions on items based on tremendous multi-field categorical data (Zhang et al., 2016), which play an indispensable role in our daily lives to help people discover information about their interests and have been adopted in a wide range of online applications, such as E-commerce, social media, news feeds, and music streaming. Researchers have developed deep-learning-based recommendation models to dig feature representations flexibly. These models have been successfully deployed across a multitude of application scenarios, thereby demonstrating their widespread adoption and effectiveness.
Motivated by the advancement of large foundation models (Kirillov et al., 2023; OpenAI, 2023; Radford et al., 2021; Rombach et al., 2022) that benefit from increasing parameters, it should have been a promising trend to scale up the recommendation model size to use the data amount fully. Yet counterintuitively, as the most-weighted component critical for performance (Qu et al., 2016; Lian et al., 2018; Wang et al., 2021), the embeddings of recommendation models are typically tuned too small such as a size of 10 (Zhu et al., 2022), and thus do not adequately capture the magnitude of data. Worsely, increasing the embedding size does not sufficiently improve the performance or even hurts the model, as shown in Figure 1a. This suggests a deficiency in the model scalability of existing architecture designs, constraining the potential upper bound for recommender systems.
@@
To figure out the reason behind it, we take a spectral analysis on the learned embedding matrices based on singular value decomposition and exhibit the normalized singular values in Figure 1b. Surprisingly, most singular values are significantly small, i.e., the learned embedding matrices are nearly low-rank, which we refer to as the embedding collapse phenomenon. With the enlarged model size, the model does not learn to capture a larger dimension of information, implying a learning process with ineffective parameter utilization, which restricts the scalability.
In this work, we study the mechanism behind the embedding collapse through empirical and theoretical analysis and shed light on the two-sided effect on model scalability of the feature interaction module, the cornerstone of recommendation models to model higher-order correlations. On the one hand, interaction with collapsed embeddings will constrain the embedding learning and, thus, in turn, aggravate the collapse issue. On the other hand, the feature interaction also plays a vital role in reducing overfitting when scaling up models, which cannot be restricted or removed.
Based on our analysis, we conclude the principle to mitigate collapse without suppressing feature interaction, so that scalable models can be approached. We propose multi-embedding as a simple yet efficient design for model scaling. Multi-embedding scales the number of independent embedding sets and incorporates embedding-set-specific interaction modules to jointly capture different patterns. Our experimental results demonstrate that multi-embedding provides scalability for extensive mainstream models and significantly mitigates embedding collapse, pointing to a methodology of breaking through the size limit of recommender systems.
@@
  • To the best of our knowledge, we are the first to point out the model scalability issue in recommender systems and discover the embedding collapse phenomenon, an urgent problem to address to enhance scalability.
  • Using empirical and theoretical analysis, we shed light on the two-sided effect of the feature interaction process on scalability based on the collapse phenomenon. We reveal that feature interaction leads to collapse while providing essential overfitting resistance.
  • Following our concluded principle to mitigate collapse without suppressing feature interaction, we propose multi-embedding as a simple unified design, consistently improving scalability and effectively mitigating embedding collapse for extensive state-of-the-art recommendation models.
@@ 2 Preliminaries
Recommendation models aim to predict an action based on features from various fields. Throughout this paper, we consider the fundamental scenario of recommender systems, in which categorial features and binary outputs are involved. Formally, suppose there are $N$ fields, with the $i$-th field denoted as $\mathcal{X}_{i}=\{1,2,...,D_{i}\}$, where $D_{i}$ denotes the field cardinality. Let
@@ 3 Embedding Collapse
Singular value decomposition has been widely used to measure the collapse phenomenon (Jing et al., 2021). In Figure 1b, we have shown that the learned embedding matrices of recommendation models are approximately low-rank with some extremely small singular values. To determine the degree of collapse for such matrices with low-rank tendencies, we propose information abundance as a generalized quantification.
Definition 3.1 (Information Abundance).
@@
Figure 2: Visualization of information abundance on the Criteo dataset. Fields are sorted by their cardinalities.
@@ 4 Feature Interaction Revisited
In this section, we delve deeper into the embedding collapse phenomenon for recommendation models. Our investigation revisits feature interaction modules which are the key to recommendation models, and revolves around two questions: (1) How is embedding collapse caused? (2) How to properly mitigate embedding collapse and enhance scalability? Through empirical and theoretical studies, we shed light on the two-sided effect of feature interaction modules on model scalability.
4.1 Interaction-Collapse Theory
To determine how feature interaction leads to embedding collapse, it is inadequate to directly analyze the raw embedding matrices since the learned embedding matrix results from interactions with all other fields, making it difficult to isolate the impact of field-pair-level interaction on embedding learning. Under this obstacle, we propose empirical evidence on models with sub-embeddings and theoretical analysis on general models, and conclude that feature interaction causes embedding collapse, named the interaction-collapse theory.
Evidence I: Empirical analysis on models with sub-embeddings.
DCNv2 (Wang et al., 2021) incorporates a crossing network parameterized with transformation matrices ${\bm{W}}_{i\to j}$ (Sun et al., 2021) over each field pair to project an embedding vector from field $i$ before interaction with field $j$. By collecting all projected embedding vectors, DCNv2 can be regarded to implicitly generate field-aware sub-embeddings ${\bm{E}}_{i}^{\to 1},{\bm{E}}_{i}^{\to 2},...,{\bm{E}}_{i}^{\to N}$ to interact with all fields from embedding matrix ${\bm{E}}_{i}$, using
@@ 4.1 Interaction-Collapse Theory
DCNv2 consists of multiple stacked cross layers, and we only discuss the first layer as simplification. To determine the collapse of sub-embedding matrices, we calculate $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ for all $i,j$ pair and show them in Figure 3a. For convenience, we pre-sort the field indices by the ascending order of information abundance, i.e., $i$ is ordered according to $\mathrm{IA}({\bm{E}}_{i})$, similar to $j$. We can observe that $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ is approximately increasing along $i$, which is trivial since ${\bm{E}}_{i}^{\to j}$ is simply a projection of ${\bm{E}}_{i}$. Interestingly, another correlation can be observed that the information abundance of sub-embeddings is co-influenced by the fields it interacts with, reflected by the increasing trend along $j$, especially for those with larger $i$. For instance, we further calculate the summation of $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ over $j$ or $i$ to study the effect of the other single variable, shown in Figure 3b and Figure 3c. The increasing trend and the correlation factor confirm the co-influence of $i$ and $j$. We further analyze the IA of FFM (Juan et al., 2016) model which also consists of sub-embeddings as DCNv2, obtaining similar observations shown in Appendx H.
(a) $\mathrm{IA}({\bm{E}}_{i}^{\to j})$.
(b) $\sum_{j=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$.
(c) $\sum_{i=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$.
Figure 3: Information abundance of sub-embedding matrices for DCNv2, with field indices sorted by information abundance of corresponding raw embedding matrices. Higher or warmer indicates larger. It is observed that $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ are co-influenced by both $\mathrm{IA}({\bm{E}}_{i})$ and $\mathrm{IA}({\bm{E}}_{j})$.
Figure 4: $\mathrm{IA}({\bm{E}}_{1})$ for toy experiments. “Small” and “Large” refers to the cardinality of $\mathcal{X}_{3}$.
Evidence II: Theoretical analysis on general recommendation models.
We now theoretically present how collapse is caused by feature interaction in general models, even without sub-embedding. For simplicity, we consider an FM-style (Rendle, 2010) feature interaction. Formally, the interaction process is defined by
@@
where $h$ is the combined feature as mentioned before. Without loss of generality, we discuss one specific row ${\bm{e}}_{1}$ of ${\bm{E}}_{1}$ and keep other embedding matrices fixed. Consider a minibatch with batch size $B$. Denote $\sigma_{i,k}$ as the $k$-th singular value of ${\bm{E}}_{i}$, and denote ${\bm{u}}_{i,k}$, ${\bm{v}}_{i,k}$ as corresponding singular vectors. We have
$\frac{\partial\mathcal{L}}{\partial{\bm{e}}_{1}}$ $=\frac{1}{B}\sum_{b=1}^{B}\frac{\partial\ell^{(b)}}{\partial h^{(b)}}\cdot\frac{\partial h^{(b)}}{\partial{\bm{e}}_{1}}=\frac{1}{B}\sum_{b=1}^{B}\frac{\partial\ell^{(b)}}{\partial h^{(b)}}\cdot\sum_{i=2}^{N}{\bm{E}}_{i}^{\top}\bm{1}_{x_{i}^{(b)}}$
@@
$=\sum_{i=2}^{N}\sum_{k=1}^{K}\alpha_{i,k}\sigma_{i,k}{\bm{v}}_{i,k}=\sum_{i=2}^{N}{\bm{\theta}}_{i},$
$\text{where}\quad{\bm{\theta}}_{i}=\sum_{k=1}^{K}\alpha_{i,k}\sigma_{i,k}{\bm{v}}_{i,k}.$
The equation means that the gradient can be decomposed into field-specific terms. We consider the component ${\bm{\theta}}_{i}$ for a certain field $i$, which is further decomposed into spectra for the corresponding embedding matrix ${\bm{E}}_{i}$. From the form of ${\bm{\theta}}_{i}$, it is observed that $\{\alpha_{i,k}\}$ are $\bm{\sigma}_{i}$-agnostic scalars determined by the training data and objective function. Thus, the variety of $\bm{\sigma}_{i}$ significantly influences the composition of ${\bm{\theta}}_{i}$. For those larger $\sigma_{i,k}$, the gradient component ${\bm{\theta}}_{i}$ will be weighted more heavily along the corresponding spectra ${\bm{v}}_{i,k}$. When ${\bm{E}}_{i}$ is low-information-abundance, the components of ${\bm{\theta}}_{i}$ weigh imbalancely, resulting in the degeneration of ${\bm{e}}_{1}$. Since different ${\bm{e}}_{1}$ affects only $\alpha_{i,k}$ instead of $\sigma_{i,k}$ and ${\bm{v}}_{i,k}$, all rows of ${\bm{E}}_{1}$ degenerates in similar manners and finally form a collapsed matrix.
To further illustrate, we conduct a toy experiment over synthetic data. Suppose there are $N=3$ fields, and we set $D_{3}$ to different values with $D_{3}<K$ and $D_{3}\gg K$ to simulate low-information-abundance and high-information-abundance cases, which matches the diverse range of the field cardinality in real-world scenarios. We train ${\bm{E}}_{1}$ while keeping ${\bm{E}}_{2},{\bm{E}}_{3}$ fixed. Details of experiment setups are discussed in Appendix G. We show the information abundance of ${\bm{E}}_{1}$ along the training process for the two cases in Figure 4. It is observed that interacting with a low-information-abundance matrix will result in a collapsed embedding.
Summary: How is collapse caused in recommendation models?
Evidence I highlights that interacting with a low-information-abundance field will result in a more collapsed sub-embedding. By considering the fact that sub-embeddings reflect the effect when fields interact since it originates from raw embeddings, we recognize the inherent mechanism of feature interaction to cause collapse, which is further confirmed by our theoretical analysis. We conclude the interaction-collapse theory:
Finding 1 (Interaction-Collapse Theory). In feature interaction of recommendation models, fields with low-information-abundance embeddings constrain the information abundance of other fields, resulting in collapsed embedding matrices.
The interaction-collapse theory generally suggests that feature interaction is the primary catalyst for collapse, thereby imposing constraints on the ideal scalability.